### v2

@@ 1 Introduction
Among cancer therapies, oncolytic virotherapy stands out as a promising avenue that harnesses the natural capabilities of viruses to selectively target and destroy cancer cells while sparing healthy tissues, [blanchette23, fountzilas17, Kelly2007651review, lawler17, russell18]. Despite its potential, the clinical efficacy of oncolytic virotherapy faces significant challenges, many of which stem from the complex dynamics of the tumour microenvironment (TME) [jin21].
@@
The TME significantly influences viral distribution and therapeutic efficacy. Factors such as extracellular matrix composition, immune cell infiltration and hypoxic regions can impede viral penetration, replication and spread within the tumour [wojton10]. These barriers create unpredictable, stochastic events that affect the consistency of viral delivery and the overall success of the therapy [jin21]. Another interesting interaction involves the neovasculature that the tumour originates through angiogenesis: most oncolytic viruses disrupt it by targeting tumour-associated vascular endothelial cells [jin21, wojton10]. On the one hand, the decrease in nutrient inflow contributes to slowing down cancer growth and, in this respect, oncolytic virotherapy could act in the same way as antiangiogenic therapy [santry20]. On the other hand, blood vessels also play a vital role in the arrival of viral particles and immune cells [santry20].
@@
The inadequate vascularisation within tumours reduces oxygen availability: this hypoxic condition is a hallmark of solid tumours, significantly affects the tumour’s evolution and profoundly impacts treatment efficacy [zhuang23]. Hypoxic regions within tumours promote aggressiveness by harbouring cells with reduced metabolic activity and heightened resistance to therapies [zhuang23]. In the context of oncolytic virotherapy, the impact of hypoxia on therapeutic efficacy is not straightforward [shengguo11]. In general, hypoxic regions within tumours harbour cells less susceptible to viral infection and replication due to reduced metabolic activity and altered cellular signalling pathways; furthermore, the physiological adaptations of tumour cells to hypoxia, such as enhanced glycolytic metabolism and resistance to apoptosis, can confer resistance to viral-induced cell death. It is important to remark that the previous considerations are not universal: indeed, some particular oncolytic viruses can specifically target receptors that are upregulated in case of the lack of oxygen [shengguo11, sadri23].
@@
Understanding how hypoxia influences the interaction between OVs and tumour cells is crucial for optimising treatment strategies and overcoming therapeutic resistance. Mathematical modelling is pivotal in unravelling these complexities and optimising therapeutic outcomes [byrne10]. The dynamics of viral replication and tumour growth have been analysed through several modelling approaches, ranging from ordinary differential equations (ODEs) [almuallem21, eftimie18, jenner19, jenner18insights, komarova10, novozhilov06, storey20, vithanage23, wodarz01] and partial differential equations (PDEs) [alzahrani19, friedman03, friedman18, kim14, pooladvand22, wu01, wu04] to stochastic agent-based models [jenner20voronoi, storey21, surendran23, wodarz12] and hybrid multi-scale models [jenner22, paiva09]. Some of these models also take into account aspects related to the tumour microenvironment, such as spatial constraints due to the extracellular matrix [alzahrani19, jenner22, kim14, pooladvand22] and interactions with the immune system [almuallem21, eftimie18, storey20, vithanage23, friedman18, jenner22, wodarz01, wu04]. Only a few mathematical models describe virotherapy in hypoxic conditions. In Ref. [ramaj23], the authors consider ODEs and nested ODEs (corresponding to infections in adjacent lymph nodes) under the assumption that oxygen concentration directly influences the infection rate. The only spatial model that we know in this context is the one presented in Ref. [boemo19], in which the difficulties of treating hypoxic regions with standard therapies motivate the use of macrophages that release oncolytic viral particles when experiencing low oxygen concentrations. The lack of spatial models that consider the influence of oxygen concentration on virotherapy effectiveness motivates the present work.
@@
Even in the absence of viral infection, non-genetic changes in cancer cells significantly affect the metabolic activity and these evolutionary dynamics can be analysed with mathematical models [bosque23]. In particular, the influence of hypoxia on tumour dynamics has been investigated through several approaches, using either discrete compartments that evolve according to different dynamics [astanin08, Gatenby, martinez12, rocha2021persistent] or a continuous spectrum of adaptation levels [ardaseva2020fluctuating, ardaseva20periodic, gallaher]. The latter approach in the continuous settings leads to integro-differential equations (IDEs) and, when spatial heterogeneity is included, to partial integro-differential equations (PIDEs) [chiari23heterogeneity, fiandaca2020mathematical, lorenzi2018role, villa21modeling]. We build upon this modelling approach and consider the heterogeneous effects of oncolytic virotherapy within the framework of spatial and phenotypic variability.
@@
A common characteristic of continuous structured models is the presence of a trade-off between different features [lorenzi24preprint]. At a cellular level, the term trade-off refers to the presence of mutually exclusive properties, usually caused by the need for metabolic optimisation, with the possibility that an improvement in one aspect necessarily leads to a deterioration in another. One of the most considered trade-offs in the context of hypoxia involves the proliferation rate and resistance to hypoxic conditions. This can be interpreted with the following meaning: cells whose proliferative performance is significantly enhanced by oxygen are also more easily induced to die in harsh conditions. In contrast, cells that are not very sensitive to increased proliferation due to the abundance of nutrients maintain a state of quiescence or low proliferation in the event of a lack of oxygen, without incurring cell death.
@@
It is important to remark that the use of a continuous trait variable to describe the system’s heterogeneity has been widely employed for equations that share some formal similarities to the ones analysed in the present work, both in the epidemiological settings [almeida21, bernardi22, lorenzi24, lorenzi21] and the ecological settings [dearaujo24, delitala13evolution]. In oncolytic virotherapy, structured populations have been used in Ref. [karev06] to model various kinds of heterogeneity, including susceptibility to infections, death rates, and virulence; neither of them considers trade-offs similar to the ones we aim to investigate. A slightly different approach involves the use of an age structure: indeed, the inclusion of the time from the infection as a structuring variable for the infected individuals dates back to the first modern SI model [kermack27]. Analogous models were also adopted in Refs. [ding22, gao22] in relation to the time from infection for oncolytic virotherapy and in Ref. [portillo24] in relation to the temporal evolution of stem cells.
As previously mentioned, in this work, we build upon the model presented in Ref. [lorenzi21] and include spatial heterogeneity and environmental factors that influence the dynamics (as in Refs. [chiari23heterogeneity, chiari23radio]); to our knowledge, both aspects constitute a novelty with respect to the existing literature.
@@ 2.1 Model description
We remark that the terms for proliferation and selective pressure follow the modelling choices adopted in Ref. [chiari23heterogeneity]. In absence of infections and epigenetic diffusion, the cell density always grows up to the same carrying capacity $K$ irrespectively of the oxygen level: indeed, $u(t,\boldsymbol{x},y)=K\delta_{\varphi(O(\boldsymbol{x}))}(y)$ is trivially a steady state (see also Eq. (A.9) in this respect). This constitutes a significant difference with respect to some previous modelling approaches of hypoxic tumours (such as the ones employed in Refs. [ardaseva2020fluctuating, ardaseva20comparative, fiandaca2020mathematical, villa21modeling]) and allows us to consider virotherapy as the only cause of cancer reduction in hypoxic conditions. In presence of epigenetic diffusion, it is reasonable to expect a smaller equilibrium value of $\rho$ due to the coexistence of subpopulations with different epigenetic characteristics and the resulting action of the term $S$.
@@
The movement follows the same law as that of uninfected cells, as there is no reason to assume that the infection has some effect on that. All the susceptible cells that undergo infection are collected in the same population. Finally, infected cells die at rate $q_{I}$. This equation does not take into account the uptake of free viruses by cancer cells: although a few models include this process [bajzer08, pooladvand21], it is usually assumed to be negligible [almuallem21, baabdulla23, jenner19, jenner18insights, storey20, vithanage23].
We assume that infected cells do not proliferate, as the virus disrupts the cellular machinery, and are not affected by environmental conditions due to their short life. Consequently, they lack all the processes the epigenetic trait affects, which motivates their absence of epigenetic structure. This approach is the same one adopted in Ref. [lorenzi21], with the only differences that we here consider a spatial structure and infections mediated by a viral population.
@@
We assume that healthy cells initially fill the tissue up to carrying capacity and their oxygen consumption is indirectly modelled through the decay at rate $q_{O}$. As the tumour grows, healthy cells are replaced by cancer cells, which consume more oxygen: the additional consumption is captured by the term $\lambda\rho(t,\boldsymbol{x})O(t,\boldsymbol{x})$. Observe that both infected and uninfected cells are responsible for oxygen consumption in the same way and the epigenetic trait of uninfected cells does not affect the consumption (as already done in Ref. [chiari23radio]).
@@ 2.2 Theoretical insights
The stationary equilibrium values of the system described in Eq. (2.9) in the spatially homogeneous case can be computed through formal asymptotic analysis. Building upon the methods employed in Refs. [lorenzipainter22, lorenzi21invasion, villa21modeling], we introduce a small parameter $\varepsilon$ and assume that $D_{y}=\varepsilon^{2}$. Furthermore, we use the time scaling $t\mapsto\frac{t}{\varepsilon}$, which allows us to study the long-time behaviour of the system. We then make for uninfected cells the real phase WKB ansatz [barles90, evans89, fleming86], as it is common in the Hamilton–Jacobi approach presented, for example, in [diekmann05, lorz11, perthame2006transport, perthame08]
@@
Furthermore, we assume that, for every $\varepsilon>0$, the initial condition $n_{\varepsilon}(0,\boldsymbol{x},y)$ is a uniformly concave function of $y$; for example, we could think that, at $t=0$, tumour cells that occupy the same position are mainly in the same phenotypic state and so the local cell phenotypic distribution $y\mapsto u_{\varepsilon}(0,\boldsymbol{x},y)$ is a sharp Gaussian-like function (meaning that $n_{\varepsilon}(0,\boldsymbol{x},y)=-(y-\varphi_{\varepsilon}(\boldsymbol{x}))^{2}$, for some appropriate function $\varphi_{\varepsilon}$). Since $y\mapsto R(y,\rho,I)$ is a concave function of $y$, we also expect $n_{\varepsilon}$ to be a strictly concave function of $y$ [mirrahimi2015asymptotic, perthame08]. We can then define
@@
$p_{M}$ maximal duplication rate $2.88\times 10^{-2}$ [h-1] [ke00]
$p_{m}$ minimal duplication rate $1.44\times 10^{-2}$ [h-1] [martinez12]
$K$ tissue carrying capacity $10^{6}$ [cells/mm3] [lodish08]
@@
$D_{y}$ cell epigenetic diffusion coefficient $5.00\times 10^{-6}$ [h-1]] [celora21]
@@
$q_{I}$ death rate of infected cells $4.17\times 10^{-2}$ [h-1] [ganly00]
$q_{v}$ virus clearance rate $1.67\times 10^{-1}$ [h-1] [mok09]
@@
$D_{v}$ virus diffusion coefficient $3.6\times 10^{-2}$ [mm2/h] [kim06]
$O_{\max}$ maximal oxygen concentration $2.16\times 10^{-3}$ $_{O_{2}}^{3}/mm$$^{3}{{}_{\text{plasma}}}$] [mckeown14, pittman11]
$O_{M}$ oxygen normoxic threshold $1.71\times 10^{-3}$ $_{O_{2}}^{3}/mm$$^{3}{{}_{\text{plasma}}}$] [mckeown14, pittman11]
$O_{m}$ oxygen hypoxic threshold $2.28\times 10^{-4}$ $_{O_{2}}^{3}/mm$$^{3}{{}_{\text{plasma}}}$] [mckeown14, pittman11]
$q_{O}$ oxygen physiological decay rate $5.60\times 10^{2}$ [h-1] estimate based on [wagner11]
@@
$D_{O}$ oxygen diffusion coefficient $3.60$ [mm2/h] [mueller-klieser84]
@@
So far, we have focused our discussion on the spatially homogeneous situation. To our knowledge, spatial dynamics in this context have not been studied analytically, not even in simpler settings (such as the one of Ref. [lorenzi21]). Nevertheless, given the forthcoming numerical simulations, it is helpful to recall some elementary facts, with the goal of elucidating several complex dynamics that we observe. If we neglect the epigenetic structure, the one-dimensional dynamics of uninfected cells in the absence of viral infection follow the equation
@@
It is well-known that there exist travelling waves solutions of this equation with speed at least $\sqrt{D_{\boldsymbol{x}}Kp/2}$ and an initial condition with compact support evolves into a wave that travels with the minimal speed [aronson80, newman80]. The addition of infection due to cell-to-cell contact originates travelling waves of the two populations (uninfected and infected), with the uninfected proliferative cells trying to escape from the infected cells [morselli23]; the addition of a viral population does not significantly change this picture. It is important to remark that an efficient infection results in a wavefront much lower than the carrying capacity, whose invasion speed is lower than $\sqrt{D_{\boldsymbol{x}}Kp/2}$.
@@
If we had $\bar{y}(t,{x})\approx\varphi(O)$ with $O$ constant in time and space, then it is clear that the minimal wave speed would be $\sqrt{D_{\boldsymbol{x}}Kp(\varphi(O))/2}$. The numerical simulations of the next section show that this is not the case, not even when the oxygen concentration is stationary and homogeneous: indeed, a rigorous study of the travelling waves should keep into account the spatial variations of $\bar{y}$ (see Ref. [lorenzi24preprint] and the references therein), which are extremely complex the case of our system and fall beyond of the scope of the present work; further complexities arise from the fact that the invasion front is lower than $K$ because of the infection and this also affects the wave speed. Nevertheless, the previous considerations provide us a broad intuition on the behaviour of the system, which is in line with the forthcoming numerical simulations.

### v3

@@ 1 Introduction
Among cancer therapies, oncolytic virotherapy stands out as a promising avenue that harnesses the natural capabilities of viruses to selectively target and destroy cancer cells while sparing healthy tissues, [14, 31, 47, 52, 78]. Despite its potential, the clinical efficacy of oncolytic virotherapy faces significant challenges, many of which stem from the complex dynamics of the tumour microenvironment (TME) [44].
@@
The TME significantly influences viral distribution and therapeutic efficacy. Factors such as extracellular matrix composition, immune cell infiltration and hypoxic regions can impede viral penetration, replication and spread within the tumour [90]. These barriers create unpredictable, stochastic events that affect the consistency of viral delivery and the overall success of the therapy [44]. Another interesting interaction involves the neovasculature that the tumour originates through angiogenesis: most oncolytic viruses disrupt it by targeting tumour-associated vascular endothelial cells [44, 90]. On the one hand, the decrease in nutrient inflow contributes to slowing down cancer growth and, in this respect, oncolytic virotherapy could act in the same way as antiangiogenic therapy [80]. On the other hand, blood vessels also play a vital role in the arrival of viral particles and immune cells [80].
@@
The inadequate vascularisation within tumours reduces oxygen availability: this hypoxic condition is a hallmark of solid tumours, significantly affects the tumour’s evolution and profoundly impacts treatment efficacy [94]. Hypoxic regions within tumours promote aggressiveness by harbouring cells with reduced metabolic activity and heightened resistance to therapies [94]. In the context of oncolytic virotherapy, the impact of hypoxia on therapeutic efficacy is not straightforward [81]. In general, hypoxic regions within tumours harbour cells less susceptible to viral infection and replication due to reduced metabolic activity and altered cellular signalling pathways; furthermore, the physiological adaptations of tumour cells to hypoxia, such as enhanced glycolytic metabolism and resistance to apoptosis, can confer resistance to viral-induced cell death. It is important to remark that the previous considerations are not universal: indeed, some particular oncolytic viruses can specifically target receptors that are upregulated in case of the lack of oxygen [81, 79].
@@
Understanding how hypoxia influences the interaction between OVs and tumour cells is crucial for optimising treatment strategies and overcoming therapeutic resistance. Mathematical modelling is pivotal in unravelling these complexities and optimising therapeutic outcomes [17]. The dynamics of viral replication and tumour growth have been analysed through several modelling approaches, ranging from ordinary differential equations (ODEs) [2, 27, 41, 43, 51, 68, 83, 86, 88] and partial differential equations (PDEs) [3, 33, 32, 50, 73, 92, 93] to stochastic agent-based models [40, 82, 84, 89] and hybrid multi-scale models [42, 69]. Some of these models also take into account aspects related to the tumour microenvironment, such as spatial constraints due to the extracellular matrix [3, 42, 50, 73] and interactions with the immune system [2, 27, 83, 86, 32, 42, 88, 93]. Only a few mathematical models describe virotherapy in hypoxic conditions. In Ref. [76], the authors consider ODEs and nested ODEs (corresponding to infections in adjacent lymph nodes) under the assumption that oxygen concentration directly influences the infection rate. The only spatial model that we know in this context is the one presented in Ref. [15], in which the difficulties of treating hypoxic regions with standard therapies motivate the use of macrophages that release oncolytic viral particles when experiencing low oxygen concentrations. The lack of spatial models that consider the influence of oxygen concentration on virotherapy effectiveness motivates the present work.
@@
Even in the absence of viral infection, non-genetic changes in cancer cells significantly affect the metabolic activity and these evolutionary dynamics can be analysed with mathematical models [16]. In particular, the influence of hypoxia on tumour dynamics has been investigated through several approaches, using either discrete compartments that evolve according to different dynamics [8, 38, 61, 77] or a continuous spectrum of adaptation levels [5, 6, 35]. The latter approach in the continuous settings leads to integro-differential equations (IDEs) and, when spatial heterogeneity is included, to partial integro-differential equations (PIDEs) [22, 29, 55, 85]. We build upon this modelling approach and consider the heterogeneous effects of oncolytic virotherapy within the framework of spatial and phenotypic variability.
@@
A common characteristic of continuous structured models is the presence of a trade-off between different features [57]. At a cellular level, the term trade-off refers to the presence of mutually exclusive properties, usually caused by the need for metabolic optimisation, with the possibility that an improvement in one aspect necessarily leads to a deterioration in another. One of the most considered trade-offs in the context of hypoxia involves the proliferation rate and resistance to hypoxic conditions. This can be interpreted with the following meaning: cells whose proliferative performance is significantly enhanced by oxygen are also more easily induced to die in harsh conditions. In contrast, cells that are not very sensitive to increased proliferation due to the abundance of nutrients maintain a state of quiescence or low proliferation in the event of a lack of oxygen, without incurring cell death.
@@
It is important to remark that the use of a continuous trait variable to describe the system’s heterogeneity has been widely employed for equations that share some formal similarities to the ones analysed in the present work, both in the epidemiological settings [1, 12, 58, 59] and the ecological settings [23, 24]. In oncolytic virotherapy, structured populations have been used in Ref. [45] to model various kinds of heterogeneity, including susceptibility to infections, death rates, and virulence; neither of them considers trade-offs similar to the ones we aim to investigate. A slightly different approach involves the use of an age structure: indeed, the inclusion of the time from the infection as a structuring variable for the infected individuals dates back to the first modern SI model [48]. Analogous models were also adopted in Refs. [26, 37] in relation to the time from infection for oncolytic virotherapy and in Ref. [75] in relation to the temporal evolution of stem cells.
As previously mentioned, in this work, we build upon the model presented in Ref. [59] and include spatial heterogeneity and environmental factors that influence the dynamics (as in Refs. [22, 21]); to our knowledge, both aspects constitute a novelty with respect to the existing literature.
@@ 2.1 Model description
We remark that the terms for proliferation and selective pressure follow the modelling choices adopted in Ref. [22]. In absence of infections and epigenetic diffusion, the cell density always grows up to the same carrying capacity $K$ irrespectively of the oxygen level: indeed, $u(t,\boldsymbol{x},y)=K\delta_{\varphi(O(\boldsymbol{x}))}(y)$ is trivially a steady state (see also Eq. (A.9) in this respect). This constitutes a significant difference with respect to some previous modelling approaches of hypoxic tumours (such as the ones employed in Refs. [5, 4, 29, 85]) and allows us to consider virotherapy as the only cause of cancer reduction in hypoxic conditions. In presence of epigenetic diffusion, it is reasonable to expect a smaller equilibrium value of $\rho$ due to the coexistence of subpopulations with different epigenetic characteristics and the resulting action of the term $S$.
@@
The movement follows the same law as that of uninfected cells, as there is no reason to assume that the infection has some effect on that. All the susceptible cells that undergo infection are collected in the same population. Finally, infected cells die at rate $q_{I}$. This equation does not take into account the uptake of free viruses by cancer cells: although a few models include this process [10, 74], it is usually assumed to be negligible [2, 9, 41, 43, 83, 86].
We assume that infected cells do not proliferate, as the virus disrupts the cellular machinery, and are not affected by environmental conditions due to their short life. Consequently, they lack all the processes the epigenetic trait affects, which motivates their absence of epigenetic structure. This approach is the same one adopted in Ref. [59], with the only differences that we here consider a spatial structure and infections mediated by a viral population.
@@
We assume that healthy cells initially fill the tissue up to carrying capacity and their oxygen consumption is indirectly modelled through the decay at rate $q_{O}$. As the tumour grows, healthy cells are replaced by cancer cells, which consume more oxygen: the additional consumption is captured by the term $\lambda\rho(t,\boldsymbol{x})O(t,\boldsymbol{x})$. Observe that both infected and uninfected cells are responsible for oxygen consumption in the same way and the epigenetic trait of uninfected cells does not affect the consumption (as already done in Ref. [21]).
@@ 2.2 Theoretical insights
The stationary equilibrium values of the system described in Eq. (2.9) in the spatially homogeneous case can be computed through formal asymptotic analysis. Building upon the methods employed in Refs. [56, 54, 85], we introduce a small parameter $\varepsilon$ and assume that $D_{y}=\varepsilon^{2}$. Furthermore, we use the time scaling $t\mapsto\frac{t}{\varepsilon}$, which allows us to study the long-time behaviour of the system. We then make for uninfected cells the real phase WKB ansatz [11, 28, 30], as it is common in the Hamilton–Jacobi approach presented, for example, in [25, 60, 71, 70]
@@
Furthermore, we assume that, for every $\varepsilon>0$, the initial condition $n_{\varepsilon}(0,\boldsymbol{x},y)$ is a uniformly concave function of $y$; for example, we could think that, at $t=0$, tumour cells that occupy the same position are mainly in the same phenotypic state and so the local cell phenotypic distribution $y\mapsto u_{\varepsilon}(0,\boldsymbol{x},y)$ is a sharp Gaussian-like function (meaning that $n_{\varepsilon}(0,\boldsymbol{x},y)=-(y-\varphi_{\varepsilon}(\boldsymbol{x}))^{2}$, for some appropriate function $\varphi_{\varepsilon}$). Since $y\mapsto R(y,\rho,I)$ is a concave function of $y$, we also expect $n_{\varepsilon}$ to be a strictly concave function of $y$ [63, 70]. We can then define
@@
$p_{M}$ maximal duplication rate $2.88\times 10^{-2}$ [h-1] [46]
$p_{m}$ minimal duplication rate $1.44\times 10^{-2}$ [h-1] [61]
$K$ tissue carrying capacity $10^{6}$ [cells/mm3] [53]
@@
$D_{y}$ cell epigenetic diffusion coefficient $5.00\times 10^{-6}$ [h-1]] [19]
@@
$q_{I}$ death rate of infected cells $4.17\times 10^{-2}$ [h-1] [36]
$q_{v}$ virus clearance rate $1.67\times 10^{-1}$ [h-1] [64]
@@
$D_{v}$ virus diffusion coefficient $3.6\times 10^{-2}$ [mm2/h] [49]
$O_{\max}$ maximal oxygen concentration $2.16\times 10^{-3}$ $_{O_{2}}^{3}/mm$$^{3}{{}_{\text{plasma}}}$] [62, 72]
$O_{M}$ oxygen normoxic threshold $1.71\times 10^{-3}$ $_{O_{2}}^{3}/mm$$^{3}{{}_{\text{plasma}}}$] [62, 72]
$O_{m}$ oxygen hypoxic threshold $2.28\times 10^{-4}$ $_{O_{2}}^{3}/mm$$^{3}{{}_{\text{plasma}}}$] [62, 72]
$q_{O}$ oxygen physiological decay rate $5.60\times 10^{2}$ [h-1] estimate based on [87]
@@
$D_{O}$ oxygen diffusion coefficient $3.60$ [mm2/h] [66]
@@
So far, we have focused our discussion on the spatially homogeneous situation. To our knowledge, spatial dynamics in this context have not been studied analytically, not even in simpler settings (such as the one of Ref. [59]). Nevertheless, given the forthcoming numerical simulations, it is helpful to recall some elementary facts, with the goal of elucidating several complex dynamics that we observe. If we neglect the epigenetic structure, the one-dimensional dynamics of uninfected cells in the absence of viral infection follow the equation
@@
It is well-known that there exist travelling waves solutions of this equation with speed at least $\sqrt{D_{\boldsymbol{x}}Kp/2}$ and an initial condition with compact support evolves into a wave that travels with the minimal speed [7, 67]. The addition of infection due to cell-to-cell contact originates travelling waves of the two populations (uninfected and infected), with the uninfected proliferative cells trying to escape from the infected cells [65]; the addition of a viral population does not significantly change this picture. It is important to remark that an efficient infection results in a wavefront much lower than the carrying capacity, whose invasion speed is lower than $\sqrt{D_{\boldsymbol{x}}Kp/2}$.
@@
If we had $\bar{y}(t,{x})\approx\varphi(O)$ with $O$ constant in time and space, then it is clear that the minimal wave speed would be $\sqrt{D_{\boldsymbol{x}}Kp(\varphi(O))/2}$. The numerical simulations of the next section show that this is not the case, not even when the oxygen concentration is stationary and homogeneous: indeed, a rigorous study of the travelling waves should keep into account the spatial variations of $\bar{y}$ (see Ref. [57] and the references therein), which are extremely complex the case of our system and fall beyond of the scope of the present work; further complexities arise from the fact that the invasion front is lower than $K$ because of the infection and this also affects the wave speed. Nevertheless, the previous considerations provide us a broad intuition on the behaviour of the system, which is in line with the forthcoming numerical simulations.